a gas jar contains 240cm3 of carbon dioxide gas. calculate :

a) number of moles of carbon dioxide gas

b) number of molecules of carbon dioxide gas

c) mass of carbon dioxide gas


1
Expert's answer
2022-05-02T05:12:03-0400

V(CO2) = 240 cm3 = 0.24 L


a) n(CO2) = V(CO2) / Vm(CO2)

n(CO2) = 0.24 L / 22.4 L/mol = 0.011 moles


b) N(CO2) = n(CO2) * NA

N(CO2) = 0.011 moles * 6.02 ×1023 mol−1 = 6.622 ×1021 molecules


c) m(CO2) = n(CO2) * Mr(CO2)

m(CO2) = 0,011 moles * 44.01 g/mol = 0.48 g